# RG evolution and effect of intermediate new-physics on $\Delta B=1$ four-fermion operators

### v1

@@ 3.1 Case 1: BNC operator scale at $10^{4}$ GeV
In Fig. 4, we present the comparison between the computed proton decay scale, $\Lambda/\sqrt{c}$, obtained for cases with no RG running (in dotted line), with RG running without including BNC interactions (in blue), and with RG running including both SM and BNC-BNV mixing (in orange), for all initial sets given in Table. 4. These results use the RG evolutions defined in Eqs. (10-13), that contribute to the proton decay channel $p\to\pi^{0}e^{+}$. The comparison clearly demonstrates that the inclusion of RG effects, particularly the BNC-BNV mixing terms, have significant effect. The numerical values are given in Table. 5.
Figure 4: Comparison of the UV scale $\Lambda/\sqrt{c}$ for the four proton-decay operators in Eqs. (5-7), contributing to the decay channel $p\to\pi^{0}e^{+}$. The dotted line represents without RG running and the two different colour bars correspond to: (i) RG running without BNC-BNV mixing interactions (in blue) and (ii) RG running including BNC-BNV operator mixing (in orange), for different initial scale values (Sets) of each BNV Wilson coefficient. Results are computed using the beta functions in Eqs. (10-13), contributing to the decay channel $p\to\pi^{0}e^{+}$ with the running scale $\mu$ from $10^{3}$ GeV to $10^{4}$ GeV.
@@
Table 5: Values of the UV scale $\Lambda/\sqrt{c}$ obtained under 4 sets of initial, considering without RG evolution ($\Lambda^{\mathrm{w/o}}/\sqrt{c}$), with RG running without including BNC effects ($\Lambda^{\mathrm{w/SM}}/\sqrt{c}$), and with RG running including both SM and BNC-BNV mixing ($\Lambda^{\mathrm{w/mix}}/\sqrt{c}$). The last two rows show the ratios $\Lambda^{\mathrm{w/SM}}/\Lambda^{\mathrm{w/o}}$ and $\Lambda^{\mathrm{w/mix}}/\Lambda^{\mathrm{w/o}}$.
@@ 3.2 Case 2: BNC operator scale at $10^{6}$ GeV
The corresponding UV scale values $\Lambda/\sqrt{c}$ are presented in Fig. 6, comparing results without RG running (in dotted line), with RG evolution in the absence of BNC-BNV mixing (in blue), and with full BNC-BNV mixing (in orange), for different initial scale values of each BNV Wilson coefficient and the numerical limits extracted for them case are listed in Table. 6.
Figure 6: Comparison of the UV scale $\Lambda/\sqrt{c}$ as in Fig. 4, where the the coefficients are RG evolved with energy scale $\mu$ from $10^{3}$ GeV to $10^{6}$ GeV.
@@
Table 6: Values of the UV scale $\Lambda/\sqrt{c}$ obtained under 4 sets of initial conditions as in Table. 5, where the running is performed as a function of the energy scale $\mu$ between $10^{3}$ GeV and $10^{6}$ GeV.
@@ 3.3 Case 3: BNC operator scale at $10^{9}$ GeV
The values of the UV scale $\Lambda/\sqrt{c}$ obtained under three conditions: without RG evolution (in dotted line), with running in scenario without BNC-BNV mixing (in blue), and with full BNC-BNV mixing (in orange) for different initial scale values of each BNV Wilson coefficient are displayed in Fig. 8. The numerical results corresponding to this case are summarised in Table. 7.
Figure 8: Comparison of the UV scale $\Lambda/\sqrt{c}$ as in Fig. 4 and Fig. 6, where the running is performed as a function of the energy scale $\mu$ between $10^{3}$ GeV and $10^{9}$ GeV.
@@
Table 7: Values of the UV scale $\Lambda/\sqrt{c}$ obtained under the 4 initial conditions as in Table. 5, where the running is performed as a function of the energy scale $\mu$ between $10^{3}$ GeV and $10^{9}$ GeV.
@@ 5 Summary
The RG evolution should, therefore, be supplemented by all new-physics that may arise at intermediate energy thresholds between these scales. In a model independent SMEFT framework, dimension-six BNC operators at below these intermediate thresholds contribute to the RG running of the proton decay Wilson coefficients. The dominant 1-loop contribution to the BNC-BNV mixing arises from operators with top quark loops. Since collider experiments do not constrain flavour violating BNC operators involving top quarks beyond $\sim 10^{4}~\text{GeV}$, this motivates a detailed study of the RG evolution of the Wilson coefficients for proton decay operators, incorporating possible mixing. This is further supported by the fact that BNV operators containing third-generation quarks are themselves not constrained, allowing their scales to be set phenomenologically. Consequently, although the mixing between two dimension-six operators is naively expected to be subleading compared to that of dimension-seven operators, such mixing can nonetheless induce significant modifications to the corresponding beta functions.
We show that BNC-BNV operator mixing, particularly involving third-generation quarks through flavour violating operators, can significantly enhance the RG running effects and thereby lower the effective new-physics scale for generating for proton decay. This highlights the importance of incorporating flavor violation structures and higher-dimensional SMEFT contributions in nucleon decay analyses. To facilitate further exploration, we provide a Python implementation that performs the RG evolution of nucleon decay Wilson coefficients up to arbitrary SMEFT scales, allowing for generic baryon number conserving extensions. The evolved coefficients are then compared with experimental bounds on nucleon decay lifetimes to constrain the scale and structure of baryon number violating interactions.
@@ Appendix A Representative 1-loop mixing calculation
In this Appendix, we present a representative 1-loop calculation illustrating the extraction of the ultraviolet-divergent part that determines the BNV-BNC mixing coefficient $a_{ij}$. The example corresponds to the diagram topology (d) shown in Fig. 2, involving the loop formed between the operators $Q^{duql}_{1331}$ and $Q^{qu1}_{1313}$ for the BNV operator $Q^{duql}_{1111}$. In this configuration, a BNV operator is inserted on one fermion line and a BNC operator on the other, with the two connected through a closed top quark loop. The same procedure applies to other operator pairs with appropriate modifications to their Dirac and colour structures, and the complete set of contributions is summarised in Table. 3.
@@ Appendix B Index of Notations
$L^{(6)}$ dimension-six part of the SMEFT Lagrangian, containing all operators suppressed by $\Lambda^{2}$.
@@
$Q_{i,prst}$ Gauge-invariant dimension-six operator in the Warsaw basis; indices $p,r,s,t$ denote fermion generations.
@@
$Q^{\text{BNC}}_{i,prst}$ Baryon Number Conserving (BNC) dimension-six operators built from SM fields.

### v2

@@ 3.1 Case 1: BNC operator scale at $10^{4}$ GeV
In Fig. 4, we present the comparison between the computed proton decay scale, $\Lambda/\sqrt{c}$, obtained for cases with no RG running (in dotted line), with RG running without including BNC interactions (in blue), and with RG running including both SM and BNC-BNV mixing (in orange), for all initial sets given in Table. 4. These results use the RG evolutions defined in Eqs. (10-13), that contribute to the proton decay channel $p\to\pi^{0}e^{+}$. The comparison clearly demonstrates that the inclusion of RG effects, particularly the BNC-BNV mixing terms, have significant effect. The numerical values are given in Table. 5. The dependence of the proton decay scale on the BNC-BNV mixing terms is summarised in Fig. 5. The plot Fig.5(a) shows the proton decay scale obtained after RG evolution without including the BNC-BNV mixing, while Fig. 5(b) shows the effect of the mixing.
Figure 4: Comparison of the UV scale $\Lambda/\sqrt{c}$ for the four proton-decay operators in Eqs. (5-7), contributing to the decay channel $p\to\pi^{0}e^{+}$ for case 1. The dotted line represents without RG running and the two different colour bars correspond to: (i) RG running without BNC-BNV mixing interactions (in blue) and (ii) RG running including BNC-BNV operator mixing (in orange), for different initial scale values (Sets) of each BNV Wilson coefficient. Results are computed using the beta functions in Eqs. (10-13), contributing to the decay channel $p\to\pi^{0}e^{+}$ with the running scale $\mu$ from $10^{3}$ GeV to $10^{4}$ GeV.
@@
Table 5: Values of the UV scale $\Lambda/\sqrt{c}$ obtained under 4 sets of initial, for case 1, considering without RG evolution ($\Lambda^{\mathrm{w/o}}/\sqrt{c}$), with RG running without including BNC effects ($\Lambda^{\mathrm{w/SM}}/\sqrt{c}$), and with RG running including both SM and BNC-BNV mixing ($\Lambda^{\mathrm{w/mix}}/\sqrt{c}$). The last two rows show the ratios $\Lambda^{\mathrm{w/SM}}/\Lambda^{\mathrm{w/o}}$ and $\Lambda^{\mathrm{w/mix}}/\Lambda^{\mathrm{w/o}}$.
(a) without BNC-BNV mixing terms
(b) with BNC-BNV mixing terms
Figure 5: The plot shows the dependence of the proton decay operator scale on the BNC-BNV parameter space for case 1. The x-axis represents the BNC operator Wilson coefficient ($C^{BNC}_{1313}$), while, the y-axis represent the BNV Wilson coefficient ($C^{BNV}_{1331}$).
@@ 3.2 Case 2: BNC operator scale at $10^{6}$ GeV
The corresponding UV scale values $\Lambda/\sqrt{c}$ are presented in Fig. 7, comparing results without RG running (in dotted line), with RG evolution in the absence of BNC-BNV mixing (in blue), and with full BNC-BNV mixing (in orange), for different initial scale values of each BNV Wilson coefficient and the numerical limits extracted for them case are listed in Table. 6. The dependence of the proton decay scale on the BNC-BNV mixing terms is summarised in Fig. 8. The plot Fig.8(a) shows the proton decay scale obtained after RG evolution without including the BNC-BNV mixing, while Fig. 8(b) shows the effect of the mixing.
Figure 7: Comparison of the UV scale $\Lambda/\sqrt{c}$ as in Fig. 4, where the the coefficients are RG evolved with energy scale $\mu$ from $10^{3}$ GeV to $10^{6}$ GeV, for case 2.
@@
Table 6: Values of the UV scale $\Lambda/\sqrt{c}$ obtained under 4 sets of initial conditions, for case 2, as in Table. 5, where the running is performed as a function of the energy scale $\mu$ between $10^{3}$ GeV and $10^{6}$ GeV.
(a) without BNC-BNV mixing terms
(b) with BNC-BNV mixing terms
Figure 8: The plot shows the dependence of the proton decay operator scale on the BNC-BNV parameter space for case 2. The x-axis represents the BNC operator Wilson coefficient ($C^{BNC}_{1313}$), while, the y-axis represent the BNV Wilson coefficient ($C^{BNV}_{1331}$).
@@ 3.3 Case 3: BNC operator scale at $10^{9}$ GeV
The values of the UV scale $\Lambda/\sqrt{c}$ obtained under three conditions: without RG evolution (in dotted line), with running in scenario without BNC-BNV mixing (in blue), and with full BNC-BNV mixing (in orange) for different initial scale values of each BNV Wilson coefficient are displayed in Fig. 10. The numerical results corresponding to this case are summarised in Table. 7. The dependence of the proton decay scale on the BNC-BNV mixing terms is summarised in Fig. 11. The plot Fig.11(a) shows the proton decay scale obtained after RG evolution without including the BNC-BNV mixing, while Fig. 11(b) shows the effect of the mixing.
Figure 10: Comparison of the UV scale $\Lambda/\sqrt{c}$ as in Fig. 4 and Fig. 7, where the running is performed as a function of the energy scale $\mu$ between $10^{3}$ GeV and $10^{9}$ GeV, for case 3.
@@
Table 7: Values of the UV scale $\Lambda/\sqrt{c}$ obtained under the 4 initial conditions, for case 3, as in Table. 5, where the running is performed as a function of the energy scale $\mu$ between $10^{3}$ GeV and $10^{9}$ GeV.
(a) without BNC-BNV mixing terms
(b) with BNC-BNV mixing terms
Figure 11: The plot shows the dependence of the proton decay operator scale on the BNC-BNV parameter space for case 3. The x-axis represents the BNC operator Wilson coefficient ($C^{BNC}_{1313}$), while, the y-axis represent the BNV Wilson coefficient ($C^{BNV}_{1331}$).
@@ 5 Summary
The RG evolution should, therefore, be supplemented by all new-physics that may arise at intermediate energy thresholds between these scales. In a model independent SMEFT framework, dimension-6 BNC operators at below these intermediate thresholds contribute to the RG running of the proton decay Wilson coefficients. The dominant 1-loop contribution to the BNC-BNV mixing arises from operators with top quark loops. Since collider experiments do not constrain flavour violating BNC operators involving top quarks beyond $\sim 10^{4}~\text{GeV}$, this motivates a detailed study of the RG evolution of the Wilson coefficients for proton decay operators, incorporating possible mixing. This is further supported by the fact that BNV operators containing third-generation quarks are themselves not constrained, allowing their scales to be set phenomenologically. Consequently, although the mixing between two dimension-6 operators is naively expected to be subleading compared to that of dimension-seven operators, such mixing can nonetheless induce significant modifications to the corresponding beta functions.
We show that BNC-BNV operator mixing, particularly involving third-generation quarks through flavour violating operators, can significantly enhance the RG running effects and thereby lower the effective new-physics scale for generating for proton decay ($\sim 10^{7}$ GeV). Correlation and the dependence of the proton decay scale on BNC-BNV Wilson coefficients is shown in Fig.5, Fig.8 and Fig.11. This highlights the importance of incorporating flavor violation structures and higher-dimensional SMEFT contributions in nucleon decay analyses. To facilitate further exploration, we provide a Python implementation that performs the RG evolution of nucleon decay Wilson coefficients up to arbitrary SMEFT scales, allowing for generic baryon number conserving extensions. The evolved coefficients are then compared with experimental bounds on nucleon decay lifetimes to constrain the scale and structure of baryon number violating interactions.
@@ Appendix A Representative 1-loop mixing calculation
In this Appendix, we present a representative 1-loop calculation illustrating the extraction of the ultraviolet-divergent part that determines the BNC-BNV mixing coefficient $a_{ij}$. The example corresponds to the diagram topology (d) shown in Fig. 2, involving the loop formed between the operators $Q^{duql}_{1331}$ and $Q^{qu1}_{1313}$ for the BNV operator $Q^{duql}_{1111}$. In this configuration, a BNV operator is inserted on one fermion line and a BNC operator on the other, with the two connected through a closed top quark loop. The same procedure applies to other operator pairs with appropriate modifications to their Dirac and colour structures, and the complete set of contributions is summarised in Table. 3.
@@ Appendix B Index of Notations
$L^{(6)}$ dimension-6 part of the SMEFT Lagrangian, containing all operators suppressed by $\Lambda^{2}$.
@@
$Q_{i,prst}$ Gauge-invariant dimension-6 operator in the Warsaw basis; indices $p,r,s,t$ denote fermion generations.
@@
$Q^{\text{BNC}}_{i,prst}$ Baryon Number Conserving (BNC) dimension-6 operators built from SM fields.